if velocity is related to displacement as v=x^2.find the acceleration in terms of displacement.
1
Expert's answer
2019-09-17T12:07:20-0400

"v=\dot{x} = x^2\\
a = \ddot{x} = \frac{dv}{dx}\frac{dx}{dt} = 2x\cdot\dot{x} =2x^3"